# The thesan-zoom project: Population III star formation continues until the end of reionization

### v1

@@ 3 The stellar metallicity distribution
We begin our investigation by showing in Fig. 2 the stellar mass - stellar metallicity relation for all central subhaloes within the target haloes at different redshifts. All haloes with stellar masses $M_{\star}$ above $10^{5.3}{\,\rm M_{\odot}}$ follow a relatively tight relation. However, for smaller masses, we observe a significant scatter down to the imposed metallicity floor, particularly at higher redshifts. The slope of the relation remains consistent over time; however, due to previous stellar feedback, the entire relation shifts to higher metallicities over time. By performing a least squares fit to the equation:
@@
Figure 3: We show a histogram of the birth mass of all stellar particles that can be found in three high-resolution haloes at redshift $z=3$ as a function of their birth redshift and metallicity normalized by the solar metallicity. All haloes show a continuous formation of Pop III stars ($\rm Z<10^{-6}\,\text{Z}_{\odot}$) until the end of the epoch of reionization ($z\approx 5-5.5$). On average, metal-rich stars can be found in larger haloes and are born at lower redshift. Fig. 10 contains the data for the remaining 11 target haloes.
@@
As a next step, we present in Fig. 3 the metallicity and age distribution of all stars found in the target haloes at redshift $z=3$. The more massive haloes show a continuous evolution of star formation and stellar metallicity, whereas the smallest one shows artefacts of a more bursty star formation history (see the companion paper Shen et al. 2025 for a more detailed discussion of the star formation history). On average, younger stars tend to be more metal-rich, and the maximum metallicity increases with halo mass. This can be attributed to the deeper gravitational potential of more massive haloes, which reduces the efficiency of supernovae blowing enriched gas out of the halo, ultimately leading to a higher star formation efficiency (e.g. Kitayama & Yoshida, 2005; Whalen et al., 2008). In our simulations, all target haloes contain a population of extremely metal-poor stars that can form until redshift $z=5$, shortly after the end of reionization. Following this period, the minimum metallicity rapidly increases to $\approx 10^{-4}\,\text{Z}_{\odot}$, which is the maximum metallicity used in other simulations to define Pop III stars. The minimum redshift for Pop III star formation would move to $z\approx 4.5$ for this less strict criterion. We note that the value of $10^{-4}\,\text{Z}_{\odot}$ aligns with the self-consistently established metallicity floor found in simulations that directly model Pop III stars (e.g. Brauer et al., 2025). Most other studies analyzing Pop III star formation at lower redshift ended before the end of reionization, e.g. at $z=7$ (Xu et al., 2016; Jaacks et al., 2019) or $z=6$ (Johnson et al., 2013) and were unable to track the end of Pop III star formation. A notable exception is Liu & Bromm (2020a), which found a continuous but decreasing Pop III star formation until $z=4$. However, this study’s lack of radiative transfer does not enable it to self-consistently model reionization, which may explain the absence of a significant decline in the Pop III star formation rate density (SFRD) while reionization progresses.
@@
For a more quantitative analysis, we present the birth mass fraction of each stellar population as a function of age in Fig. 4. More massive haloes typically exist in over-dense regions with earlier structure formation, which leads to an earlier onset of star formation (Regan, 2023). This trend can also be seen in our simulations. Star formation begins around redshift $z\approx 24$ ($\rm 136\,Myr$), which is slightly later than the canonical value of $z\approx 30$ for the start of Pop III star formation (Klessen & Glover, 2023). This discrepancy can be explained by our lower resolution, which is used to simulate the most massive target haloes and does not allow us to fully resolve the earliest minihaloes capable of forming stars. Feedback from the first stars pollutes their birth halo with metals and leads to the formation of the first enriched Pop II stars at $z=21$ ($\rm 168\,Myr$), similar to $z\approx 22$ as found in Brauer et al. (2025). Pop II star formation starts to dominate in most target haloes around redshift 12 to 18, with an earlier transition in the more massive haloes due to the higher star formation rates. In the more massive halo, we also observe a transition to Pop I star formation at $z\approx 6-7$. Nonetheless, all haloes show a consistent formation of Pop III stars down to redshift $5$. The timing of the transition to Pop II dominated star formation is consistent with small box simulations (e.g. $z\approx 17$, Sarmento et al., 2018; Brauer et al., 2025) and large box simulations (e.g. $z\approx 13$, Venditti et al., 2023). However, the exact timing also depends on the critical transition metallicity. To quantitatively compare with previous studies, we would need global star formation rate densities for different stellar populations. In our case, these can be computed in the high-resolution regions; however, those regions are biased and would exhibit a similar spread in transition times as the target haloes. Estimates for global star formation rates can be obtained by a detailed comparison with the same regions in the parent box, as Kannan et al. (2025) did for other properties, such as the galaxy stellar mass function or the UV luminosity function. This analysis will be included in an upcoming paper.
@@ 4.1 The evolution of the gas phase
To better understand the birthplaces of Pop III stars, we analyse the high-resolution gas outside of the target halo, specifically within a comoving distance of three times the virial radius of the target halo at redshift $z=3$. For each low metallicity gas cell ($Z<10^{-6}\,\text{Z}_{\odot}$), we check the mass of its parent halo and define the intergalactic medium (IGM) as all gas outside of any halo. For the gas that is not part of the IGM and the target halo, we introduce several bins based on the mass of their parent halo. In Fig. 7, we show the temporal evolution of the mass of all low-metallicity gas, as well as the mass of high-density low-metallicity gas, which we define by using the same density threshold $n_{\mathrm{H}}=10~{}\mathrm{cm}^{-3}$ as used for star formation. This gas represents a superset of the gas which is allowed to form stars since the latter one is additionally Jeans unstable. Most of the low-metallicity gas in all simulations can be found in the IGM and low-mass haloes with masses below $10^{7}{\,\rm M_{\odot}}$. This gas continues to exist until the end of our simulations at redshift $z=3$, but its amount decreases over time due to metal enrichment. The presence of large reservoirs of pristine gas in minihaloes and the IGM is consistent with previous studies; e.g. Wise et al. (2012) found almost pristine haloes below $M_{\rm halo}\approx 10^{7.3}{\,\rm M_{\odot}}$ at $z=7$ and Liu & Bromm (2020b) found a volume filling fraction of only 2% at $z=4$ of enriched gas with $Z>10^{-4}\,\text{Z}_{\odot}$ in a cosmological box. However, detailed quantitative comparisons would require full cosmological boxes, as the regions around the target haloes can be biased. In our simulations, the pristine gas in the IGM does not reach the densities necessary for star formation; instead, Pop III star formation is dominated by haloes below $10^{9}{\,\rm M_{\odot}}$. In haloes with masses below $10^{7}{\,\rm M_{\odot}}$, a smaller fraction of low-metallicity gas reaches higher densities, which can be explained by the lower efficiency of molecular hydrogen cooling in very small haloes (Glover, 2012).
@@ 5.1 Comparison with theoretical studies
Theoretical studies of Pop III star formation can be categorized into two main categories. The first category focuses on the detailed formation processes of these stars during the period when they dominated the cosmic star formation rate (typically at $z>11$). These studies examine the internal properties of minihaloes and the conditions necessary for star formation (e.g. Smith et al., 2015; Kulkarni et al., 2021; Liu et al., 2022; Kiyuna et al., 2023; Lenoble et al., 2024; Correa Magnus et al., 2024; Smith et al., 2024). Some studies in this category also attempt to resolve an IMF using a zoom-in approach (Hirano et al., 2015; Stacy et al., 2016). The second category of studies aims to understand the statistical properties of Pop III star formation using cosmological boxes. Our study falls into the second category, but we use a zoom-in approach that allows for higher resolution, although this choice prevents us from directly obtaining the cosmic-averaged star formation rate density.
@@ 5.3 Caveats
Additionally, our chemical network does not include deuterium, particularly molecular HD. These molecules have a non-zero dipole moment and allow cooling below $\approx 100$ K (Nagakura & Omukai, 2005). It requires external Lyman Werner radiation or a massive halo (Nishijima et al., 2024) to be efficient. The reduced temperature would lower the Jeans length, potentially leading to earlier star formation in our model. As we have shown in Fig. 8, Pop III stars form at $z<8$ in subhaloes with masses below $10^{8}{\,\rm M_{\odot}}$, which we expect to be minimally affected by HD cooling. Nonetheless, HD cooling could be significant in more massive haloes at higher redshifts.
@@
Lyman-Werner radiation can travel through the neutral IGM due to its low opacity below 13.6eV (Haiman et al., 2000). As a consequence, shortly after the formation of the first Pop III stars, an extragalactic LW background forms (Incatasciato et al., 2023), which can dissociate molecular hydrogen in minihaloes, preventing them from cooling (Haiman et al., 1997). Atomic cooling haloes with $M_{\rm halo}>10^{7}{\,\rm M_{\odot}}$ are less sensitive to the LW background, as they have additional cooling channels involving atomic hydrogen, which allows them to reach higher initial densities. Only a powerful local source would be able to prevent them entirely from molecular hydrogen cooling (e.g. Regan et al., 2017). Our simulations include local LW sources self-consistently, and most of the Pop III star formation happens in haloes with $M_{\rm halo}>10^{7}{\,\rm M_{\odot}}$. Nevertheless, we stress that we only inject radiation in the UV bins as a boundary condition and no LW radiation. Including an LW background would slightly reduce the Pop III star formation rate, though it might be shifted to a lower redshift as haloes surpass the atomic cooling limit.
@@ Acknowledgements
Appendix A Additional plots per halo

### v2

@@ 3 The stellar metallicity distribution
We begin our investigation by showing in Fig. 2 the stellar mass - stellar metallicity relation for all central subhaloes within the target haloes at different redshifts. The stellar mass is measured within twice the stellar half-mass radius, and the stellar metallicity is defined as the mass-weighted mean metallicity of all star particles within the stellar half-mass radius. All haloes with stellar masses $M_{\star}$ above $10^{5.3}{\,\rm M_{\odot}}$ follow a relatively tight relation. However, for smaller masses, we observe a significant scatter down to the imposed metallicity floor, particularly at higher redshifts. The slope of the relation remains consistent over time; however, due to previous stellar feedback, the entire relation shifts to higher metallicities over time. By performing a least squares fit to the equation:
@@
Figure 3: We show a histogram of the birth mass of all stellar particles that can be found in three high-resolution haloes at redshift $z=3$ as a function of their birth redshift and metallicity normalized by the solar metallicity. All haloes contain Pop III stars ($\rm Z<10^{-6}\,\text{Z}_{\odot}$) born at the end of the epoch of reionization ($z\approx 5-5.5$), which mostly form in satellite galaxies or ex-situ. On average, metal-rich stars can be found in larger haloes and are born at lower redshift. Fig. 14 contains the data for the remaining 11 target haloes.
@@
As a next step, we present in Fig. 3 the metallicity and birth redshift distribution of all stars found in the target haloes at redshift $z=3$. The more massive haloes show a continuous evolution of star formation and stellar metallicity, whereas the smallest one shows artefacts of a more bursty star formation history (see the companion paper Shen et al. (2025) for a more detailed discussion of the star formation history). On average, younger stars tend to be more metal-rich, and the maximum metallicity increases with halo mass. This can be attributed to the deeper gravitational potential of more massive haloes, which reduces the efficiency of supernovae blowing enriched gas out of the halo, ultimately leading to a higher star formation efficiency (e.g. Kitayama & Yoshida, 2005; Whalen et al., 2008). In our simulations, all target haloes contain a population of extremely metal-poor stars that can form until redshift $z=5$, shortly after the end of reionization. Following this period, the minimum metallicity rapidly increases to $\approx 10^{-4}\,\text{Z}_{\odot}$, which is the maximum metallicity used in other simulations to define Pop III stars. The minimum redshift for Pop III star formation would move to $z\approx 4.5$ for this less strict criterion. We note that the value of $10^{-4}\,\text{Z}_{\odot}$ aligns with the self-consistently established metallicity floor found in simulations that directly model Pop III stars (e.g. Brauer et al., 2025). Most other studies analyzing Pop III star formation at lower redshift ended before the end of reionization, e.g. at $z=7$ (Xu et al., 2016; Jaacks et al., 2019) or $z=6$ (Johnson et al., 2013) and were unable to track the end of Pop III star formation. A notable exception is Liu & Bromm (2020a), which found a continuous but decreasing Pop III star formation until $z=4$. However, this study’s lack of radiative transfer does not enable it to self-consistently model reionization, which may explain the absence of a significant decline in the Pop III star formation rate density (SFRD) while reionization progresses.
@@
For a more quantitative analysis, we present the birth mass fraction of each stellar population as a function of age in Fig. 4. More massive haloes typically exist in over-dense regions with earlier structure formation, which leads to an earlier onset of star formation (Regan, 2023). This trend can also be seen in our simulations. Star formation begins around redshift $z\approx 24$ ($\rm 136\,Myr$), which is slightly later than the canonical value of $z\approx 30$ for the start of Pop III star formation (Klessen & Glover, 2023). This discrepancy can be explained by our lower resolution, which is used to simulate the most massive target haloes and does not allow us to fully resolve the earliest minihaloes capable of forming stars. Feedback from the first stars pollutes their birth halo with metals and leads to the formation of the first enriched Pop II stars at $z=21$ ($\rm 168\,Myr$), similar to $z\approx 22$ as found in Brauer et al. (2025). Pop II star formation starts to dominate in most target haloes around redshift 12 to 18, with an earlier transition in the more massive haloes due to the higher star formation rates. In the more massive halo, we also observe a transition to Pop I star formation at $z\approx 6-7$. Nonetheless, in all three zoom-in regions, we find the formation of Pop III stars down to redshift $5$. The timing of the transition to Pop II dominated star formation is consistent with small box simulations (e.g. $z\approx 17$, Sarmento et al., 2018; Brauer et al., 2025), large box simulations (e.g. $z\approx 13$, Venditti et al., 2023) and semi-analytic models (e.g. $z\approx 15$, Hartwig et al., 2024). However, the exact timing also depends on the critical transition metallicity. To quantitatively compare with previous studies, we would need global star formation rate densities for different stellar populations. In our case, these can be computed in the high-resolution regions; however, those regions are biased and would exhibit a similar spread in transition times as the target haloes. Estimates for global star formation rates can be obtained by a detailed comparison with the same regions in the parent box, as Kannan et al. (2025) did for other properties, such as the galaxy stellar mass function or the UV luminosity function. This analysis will be included in an upcoming paper.
@@ 4.1 The evolution of the gas phase
To better understand the birthplaces of Pop III stars, we analyse the high-resolution gas outside of the target halo, specifically within a comoving distance of three times the virial radius of the target halo at redshift $z=3$. For each low metallicity gas cell ($Z<10^{-6}\,\text{Z}_{\odot}$), we check the mass of its parent halo and define the intergalactic medium (IGM) as all gas outside of any halo. For the gas that is not part of the IGM and the target halo, we introduce several bins based on the mass of their parent halo. In Fig. 7, we show the temporal evolution of the mass of all low-metallicity gas, as well as the mass of high-density low-metallicity gas, which we define by using the same density threshold $n_{\mathrm{H}}=10~{}\mathrm{cm}^{-3}$ as used for star formation. This gas represents a superset of the gas which is allowed to form stars since the latter one is additionally Jeans unstable. Most of the low-metallicity gas in all simulations can be found in the IGM and low-mass haloes with masses below $10^{7}{\,\rm M_{\odot}}$. This gas continues to exist until the end of our simulations at redshift $z=3$, but its amount decreases over time due to metal enrichment. Nevertheless, metal mixing in the IGM seems to be inefficient. The presence of large reservoirs of pristine gas in minihaloes and the IGM is consistent with previous studies; e.g. Wise et al. (2012) found almost pristine haloes below $M_{\rm halo}\approx 10^{7.3}{\,\rm M_{\odot}}$ at $z=7$ and Liu & Bromm (2020b) found a volume filling fraction of only 2% at $z=4$ of enriched gas with $Z>10^{-4}\,\text{Z}_{\odot}$ in a cosmological box. However, detailed quantitative comparisons would require full cosmological boxes, as the regions around the target haloes can be biased. In our simulations, the pristine gas in the IGM does not reach the densities necessary for star formation; instead, Pop III star formation is dominated by haloes below $10^{9}{\,\rm M_{\odot}}$. In haloes with masses below $10^{7}{\,\rm M_{\odot}}$, a smaller fraction of low-metallicity gas reaches higher densities, which can be explained by the lower efficiency of molecular hydrogen cooling in very small haloes (Glover, 2012).
@@ 5.1 Comparison with theoretical studies
Theoretical studies of Pop III star formation can be categorized into two main categories. The first category focuses on the detailed formation processes of these stars during the period when they dominated the cosmic star formation rate (typically at $z>11$). These studies examine the internal properties of minihaloes and the conditions necessary for star formation (e.g. Abel et al., 2002; Yoshida et al., 2003; O’Shea & Norman, 2007; Smith et al., 2015; Kulkarni et al., 2021; Liu et al., 2022; Kiyuna et al., 2023; Lenoble et al., 2024; Correa Magnus et al., 2024; Smith et al., 2024). Some studies in this category also attempt to resolve an IMF using a zoom-in approach (Hirano et al., 2015; Stacy et al., 2016). The second category of studies aims to understand the statistical properties of Pop III star formation using cosmological boxes. Our study falls into the second category, but we use a zoom-in approach that allows for higher resolution, although this choice prevents us from directly obtaining the cosmic-averaged star formation rate density.
@@ 5.3 Caveats
Additionally, our chemical network does not include deuterium, particularly molecular HD. HD formation becomes efficient below $200$ K due to its lower binding energy compared to molecular hydrogen (chemical fractionation). These molecules have a non-zero dipole moment and allow cooling below $\approx 100$ K down to the redshift-dependent CMB temperature floor (Nagakura & Omukai, 2005). The low temperatures required for efficient HD formation can be achieved in various ways—for example, in haloes exposed to external Lyman-Werner radiation (Nishijima et al., 2024), in more massive haloes, or simply through a slower gravitational collapse (Greif et al., 2011b). The reduced temperature would lower the Jeans length, potentially leading to earlier star formation in our model. As we have shown in Fig. 8, Pop III stars form at $z<8$ in subhaloes with masses below $10^{8}{\,\rm M_{\odot}}$, where we expect the impact of HD cooling to be minimal. Moreover, HD formation itself requires cold temperatures and high densities—conditions that, in our model, already lead to star formation.
@@
As discussed in more detail in appendix C, our chemical network assumes that all $\mathrm{H}^{-}$ ions are converted into molecular hydrogen, thereby neglecting destruction by photodetachment and mutual neutralization. As shown in Fig. 13, this is generally a good approximation, but may break down in dense gas near newly formed Pop III stars. Nevertheless, since a Pop III star has already formed in these cases, a more comprehensive chemical network would likely result only in a slight reduction in the overall Pop III formation rate.
Lyman-Werner radiation can travel through the neutral IGM due to its low opacity below 13.6eV (Haiman et al., 2000). As a consequence, shortly after the formation of the first Pop III stars, an extragalactic LW background forms (Incatasciato et al., 2023), which can dissociate molecular hydrogen in minihaloes, preventing them from cooling (Haiman et al., 1997). Atomic cooling haloes with $M_{\rm halo}>10^{7}{\,\rm M_{\odot}}$ are less sensitive to the LW background, as they have additional cooling channels involving atomic hydrogen, which allows them to reach higher initial densities. Only a powerful local source would be able to prevent them entirely from molecular hydrogen cooling (e.g. Regan et al., 2017). Our simulations include local LW sources self-consistently, and most of the Pop III star formation happens in haloes with $M_{\rm halo}>10^{7}{\,\rm M_{\odot}}$. Nevertheless, we stress that we only inject radiation in the UV bins as a boundary condition and no LW radiation. To evaluate the strength of local LW sources, we calculate in appendix B the mean and median LW intensity in the zoom-in regions as a function of redshift. In zoom-in regions with more massive halos, the local radiation field exceeds the expected cosmic background, suggesting that adding the latter one would, in these simulations, probably not stop Pop III star formation. Nevertheless, including an LW background could slightly reduce the Pop III star formation rate in simulations with only a small target halo, though it might be shifted to a lower redshift as haloes surpass the atomic cooling limit.
@@ Acknowledgements
Appendix A The density and temperature distribution of star-forming gas
Figure 10: We present the mass-weighted, normalized probability density of the gas density within Voronoi cells immediately prior to the formation of a star particle. Results are shown for three different resolution levels, with higher-resolution cases (8× and 16×) limited to zoom-in simulations of 9 and 5 regions, respectively. The distributions are further separated by stellar metallicity, distinguishing Pop I/II stars (red) from Pop III stars (blue). The distribution shows only a weak dependence on redshift but systematically shifts toward higher densities with increasing resolution. Pop III stars tend to form in denser gas and exhibit a narrower distribution, reflecting the typically higher temperatures and narrower phase structure of their birth environment.
Figure 11: Same as Fig. 10, but showing the gas temperature at the time of star formation instead of density. Pop III stars form from warmer gas compared to Pop I/II stars. As the resolution increases, the temperature at star formation decreases, consistent with higher gas densities and more efficient cooling prior to star formation.
As discussed in section 2.1, gas cells are eligible to form stars in thesan-zoom when they become Jeans unstable—that is, when their effective cell size falls below the local Jeans length. Additionally, we impose a density threshold of $n_{\mathrm{H}}=10~{}\mathrm{cm}^{-3}$, to prevent star formation in cold, low-density cells. Within the ISM, we expect the Jeans instability criterion to be the more restrictive of the two. Therefore, in Fig. 10, we present the density distribution of Voronoi cells immediately prior to the formation of star particles, considering all high-resolution stars in the full zoom-in regions. All Pop III stars form in gas that is significantly denser than the imposed threshold. As the resolution increases, the typical star formation density also shifts to higher values, which directly reflects the ability to resolve smaller Jeans lengths. Pop III stars typically form in denser gas than Pop I/II stars, as the absence of efficient metal-line cooling leads to higher temperatures in their star-forming regions. At the highest resolution, the gas density distribution for Pop I/II star-forming cells exhibits a double-peaked structure, with the higher-density peak associated with low-metallicity gas. In post-processing, we also calculate the temperature distribution of star-forming cells, assuming the birth mass of each star particle corresponds to the Jeans mass of the original gas cell. This assumption is justified by our adoption of a high star formation efficiency. As shown in Fig. 11, Pop III stars indeed form in warmer gas, with a lower temperature limit of approximately 200 K—consistent with the minimum temperature achievable through molecular hydrogen cooling. At higher resolutions, we are able to resolve smaller collapsing structures, which in turn allows us to capture lower temperatures during star formation.
Appendix B The local Lyman-Werner background
Figure 12: The temporal evolution of the Lyman-Werner radiation intensity in units of $J_{21}=10^{-21}\mathrm{erg\,s^{-1}\,Hz^{-1}\,sr^{-1}\,cm^{-2}}$ in the high-resolution IGM for all 14 zoom-in regions. We present the median value (blue line), the volume-weighted average (green line), and the fitting function from Incatasciato et al. (2023), which post-processed the FiBY simulation suite and is valid for $z>6$. In all simulations, local sources gradually build up an LW background, with this process beginning earlier in zoom-in regions that host larger target haloes.
As discussed in section 5, we do not include an external Lyman-Werner (LW) radiation background, considering only local sources. To assess whether these local sources are sufficient to generate an LW background, we calculate the time evolution of the median and volume-weighted mean LW radiation intensity in each zoom-in region, focusing on the high-resolution intergalactic medium (IGM). Here, we define the IGM as all gas cells not associated with any halo. Consistent with previous studies, we express the intensity in units of $J_{21}=10^{-21}\mathrm{erg\,s^{-1}\,Hz^{-1}\,sr^{-1}\,cm^{-2}}$, which is typically calculated at the Lyman limit. For this calculation, we assume a flat spectrum across the LW band and compare our results, shown in Fig. 12, with the fit from Incatasciato et al. (2023), which is valid for $6\leq z\leq 23$. As expected, the LW background is stronger in zoom-in regions with more massive target halos. In our representative examples, m12.6_4x and m10.8_8x, the mean LW radiation intensity is comparable to or exceeds the predicted cosmological LW background. In contrast, simulations with the smallest target halos show suppressed star formation after reionization, leading to a correspondingly weaker LW background due to the scarcity of local sources. We note that this calculation was performed using the full zoom-in region, so we expect a higher LW background on average close to the local sources.
Appendix C The formation of molecular hydrogen in thesan-zoom
Figure 13: The destruction rates of $\mathrm{H}^{-}$ by associative detachment (AD), photodetachment (PD), and mutual neutralization (MN) as a function of density in primordial gas. We use simulations with zoom factor 16 and consider gas with metallicity below $10^{-6}Z_{\odot}$. The solid lines represent the median values, while the shaded regions indicate the 16th–84th and 2nd–98th percentile intervals, respectively. Molecular hydrogen formation through AD typically dominates; however, photodetachment can become significant in regions near newly formed stars.
Coefficient Value Source
$\alpha_{\mathrm{H}_{2}}^{D}$
$\frac{9.0\times 10^{-17}\,T_{2}^{0.5}}{1+0.4T_{2}^{0.5}+0.2T_{2}+0.08T_{2}^{2}}\,\text{cm}^{3}\,\text{s}^{-1}$
Nickerson et al. (2018)
$\alpha_{\mathrm{H}_{2}}^{\mathrm{GP}}$
$1.83\times 10^{-18}\,T_{K}^{0.88}\,\text{cm}^{3}\,\text{s}^{-1}$
Hutchins (1976); McKee & Krumholz (2010)
$\alpha_{\mathrm{H}_{2}}^{\mathrm{3B}}$
$6\times 10^{-32}T_{K}^{-0.25}+2\times 10^{-31}T_{K}^{-0.5}\,\text{cm}^{6}\,\text{s}^{-1}$
Forrey (2013); Nickerson et al. (2018)
$\sigma_{\mathrm{H}_{2}\mathrm{H\,I}}$
$7.073\times 10^{-19}\,T_{K}^{2.012}\cdot\frac{e^{-5.179\times 10^{4}/T_{K}}}{\left(1+2.130\times 10^{-5}T_{K}\right)^{3.512}}\,\text{cm}^{3}\,\text{s}^{-1}$
Dove & Mandy (1986); Nickerson et al. (2018)
$\sigma_{\mathrm{H}_{2}\mathrm{H}_{2}}$
$5.996\times 10^{-30}\,T_{K}^{4.1881}\cdot\frac{e^{-5.466\times 10^{4}/T_{K}}}{\left(1+6.761\times 10^{-6}T_{K}\right)^{5.6881}}\,\text{cm}^{3}\,\text{s}^{-1}$
Martin et al. (1998); Nickerson et al. (2018)
Table 3: The coefficients we use in the chemical network for molecular hydrogen chemistry. The implementation is further discussed in Kannan et al. (2020), which itself is based on Nickerson et al. (2018). $T_{K}$ is here the temperature in Kelvin and $T_{2}=T_{K}/100$
thesan-zoom follows the non-equilibrium abundances of H i, H ii, He i, He ii, He iii and H 2. The reactions governing the ionization and recombination of atomic hydrogen and helium are described in detail in the Appendix B of Kannan et al. (2019), to which we refer for further information. The evolution of the molecular hydrogen abundance is given by:
$\frac{dn_{\mathrm{H}_{2}}}{dt}=$ $\alpha_{\mathrm{H}_{2}}^{D}\left(\frac{D}{D_{\mathrm{MW}}}\right)n_{\mathrm{H}}n_{\mathrm{H\,I}}+\alpha_{\mathrm{H}_{2}}^{\mathrm{GP}}n_{\mathrm{H\,I}}n_{e}$ (2)
$+\alpha_{\mathrm{H}_{2}}^{\mathrm{3B}}n_{\mathrm{H\,I}}^{2}\left(n_{\mathrm{H\,I}}+\frac{n_{\mathrm{H}_{2}}}{8}\right)-\sigma_{\mathrm{H}_{2}\mathrm{H\,I}}n_{\mathrm{H}_{2}}n_{\mathrm{H\,I}}$ (3)
$-\sigma_{\mathrm{H}_{2}\mathrm{H}_{2}}n_{\mathrm{H}_{2}}^{2}-n_{\mathrm{H}_{2}}\left(S_{\mathrm{H}_{2}}\Gamma_{\mathrm{H}_{2}}^{\mathrm{LW}}+\Gamma_{\mathrm{H}_{2}}^{+}\right),$ (4)
where $\alpha_{\mathrm{H}_{2}}^{D}$ is the rate of H2 formation on dust grains, with $D$ and $D_{\mathrm{MW}}=0.01$ denoting the self-consistently evolved local and Milky Way dust-to-gas ratios, respectively. The term $\alpha_{\mathrm{H}_{2}}^{\mathrm{GP}}$ accounts for low density gas-phase formation, while $\alpha_{\mathrm{H}_{2}}^{\mathrm{3B}}$ describes the formation through three-body processes, which become efficient at high densities. The destruction of molecular hydrogen occurs via collisions with atomic and molecular hydrogen, represented by the rates $\sigma_{\mathrm{H}_{2}\mathrm{H\,I}}$ and $\sigma_{\mathrm{H}_{2}\mathrm{H}_{2}}$, respectively. The final term describes photodissociation and photoionization. $\Gamma_{\mathrm{H}_{2}}^{\mathrm{LW}}$ is the local Lyman–Werner band flux responsible for photodissociation, modulated by the self-shielding factor $S_{\mathrm{H}_{2}}$. The self-shielding approximation we adopt follows equation (4) of Kannan et al. (2020), based on Draine & Bertoldi (1996), as our radiative transfer scheme does not capture self-shielding by line overlap effects. The final term $\Gamma_{\mathrm{H}_{2}}^{+}$ accounts for photoionization by photons with energies above 15.2eV. The coefficients used in our chemical model are summarized in Table 3. In dust-free environments, molecular hydrogen formation proceeds exclusively via gas-phase reactions, with the three-body channel becoming efficient only at high densities $n_{\mathrm{H}}>10^{9}\,\text{cm}^{-3}$ (Klessen & Glover, 2023). In low-density gas molecular hydrogen can form using $\mathrm{H}^{-}$ or $\mathrm{H}^{+}$, with the first one being the dominant one except for rare cases when there are more free protons than electrons or when $\mathrm{H}^{-}$ photodetachment is much more effective than $\mathrm{H}^{+}$ photodissociation:
[…]
Radiative attachment exhibits a significantly lower reaction rate compared to associative detachment. Therefore, we adopt the reaction rate of radiative attachment as the effective formation rate of molecular hydrogen (similar to McKee & Krumholz, 2010). Rather than explicitly tracking the abundance of $\mathrm{H}^{-}$, we assume it is in equilibrium between its formation via radiative attachment and its destruction via associative detachment. This implies that all $\mathrm{H}^{-}$ is efficiently converted into molecular hydrogen, and that alternative destruction pathways are subdominant. This assumption is valid if associative detachment dominates over other destruction mechanisms, which generally include:
[…]
To assess the relative importance of these three destruction channels, we adopt the following reaction rates:
$R_{\mathrm{AD}}$ $\approx\left(1.3\times 10^{-9}\right)\cdot n_{\mathrm{H}}\cdot n_{\mathrm{H}^{-}}\quad\text{cm}^{-3}\,\text{s}^{-1}$ (9)
$R_{\mathrm{MN}}$ $=\left(7\times 10^{-7}\cdot T^{-0.5}\right)\cdot n_{\mathrm{H}^{+}}\cdot n_{\mathrm{H}^{-}}\quad\text{cm}^{-3}\,\text{s}^{-1}$ (10)
$R_{\mathrm{PD}}$ $=\left(8.8\times 10^{-10}\cdot J_{21}\right)\cdot n_{\mathrm{H}^{-}}\quad\text{s}^{-1},$ (11)
using rate coefficients from McKee & Krumholz (2010), Dalgarno & Lepp (1987); Glover et al. (2006) and Omukai (2001), respectively. We note that the photodetachment rate is particularly sensitive to the assumed shape of the radiation spectrum (Agarwal & Khochfar, 2015). Assuming charge neutrality, we approximate $n_{\mathrm{H}^{+}}\approx n_{e}$. We compute all $\mathrm{H}^{-}$ destruction rates for high-resolution gas with metallicity $Z<10^{-6}Z_{\odot}$, and present the median values as a function of hydrogen number density in Fig. 13, using our highest-resolution simulations. The spread in each rate is shown via the 16th–84th and 2nd–98th percentile intervals. The rate of associative detachment scales linearly with density and thus exhibits no intrinsic scatter. In contrast, the photodetachment rate depends on the local radiation field. Its median value decreases with increasing density due to enhanced shielding, but rises sharply at $n_{\mathrm{H}}>10^{5}\,\mathrm{cm}^{-3}$, coinciding with the onset of Pop III star formation in nearby cells (see also Fig. 10). Photodetachment can dominate $\mathrm{H}^{-}$ destruction in low-density gas and even in denser regions ($n_{\mathrm{H}}\gtrsim 200\,\mathrm{cm}^{-3}$) where local Lyman–Werner radiation from newly formed Pop III stars becomes significant. Mutual neutralization depends on the local free electron density and temperature, and is more efficient at lower temperatures. However, it remains subdominant compared to associative detachment due to the low ionization fractions in our simulations, typically less than $10^{-3}$ for gas with $n_{\mathrm{H}}>1\,\mathrm{cm}^{-3}$. These results are consistent with our lower-resolution simulations, although star formation occurs at slightly lower densities, shifting the density region in which photodetachment becomes the dominant destruction mechanism. Overall, molecular hydrogen formation remains inefficient in primordial gas, with molecular fractions rarely exceeding 1%. The median value is typically around $10^{-3}$ for gas denser than $100\,\mathrm{cm}^{-3}$.
Appendix D Additional plots per halo